# Predictively Oriented Posteriors

### v1

@@ Abstract
We advocate for a new statistical principle that combines the most desirable aspects of both parameter inference and density estimation. This leads us to the predictively oriented (PrO) posterior, which expresses uncertainty as a consequence of predictive ability. Doing so leads to inferences which predictively dominate both classical and generalised Bayes posterior predictive distributions: up to logarithmic factors, PrO posteriors converge to the predictively optimal model average at rate $n^{-1/2}$. Whereas classical and generalised Bayes posteriors only achieve this rate if the model can recover the data-generating process, PrO posteriors adapt to the level of model misspecification. This means that they concentrate around the true model at rate $n^{1/2}$ in the same way as Bayes and Gibbs posteriors if the model can recover the data-generating distribution, but do not concentrate in the presence of non-trivial forms of model misspecification. Instead, they stabilise towards a predictively optimal posterior whose degree of irreducible uncertainty admits an interpretation as the degree of model misspecification—a sharp contrast to how Bayesian uncertainty and its existing extensions behave. Lastly, we show that PrO posteriors can be sampled from by evolving particles based on mean field Langevin dynamics, and verify the practical significance of our theoretical developments on a number of numerical examples.
@@
In the context of prediction however, posterior collapse is generally not desirable: it implies that as the sample size increases, marginalising over the Bayes posterior ceases to function as a genuine combination of individual predictive distributions. Instead of averaging over a model class, such predictive distributions essentially predict from a single model. Crucially, this behaviour persists even for generalised Bayesian methods that replace a likelihood-based approach with a loss-based belief update (Bissiri et al., 2016; Jiang and Tanner, 2008; Zhang, 2006; Knoblauch et al., 2022; Alquier, 2024; Syring and Martin, 2019; Martin and Syring, 2022) As a result, the posterior predictives obtained from both classical and generalised Bayes approaches quickly resemble those obtained from a point estimator as sample size grows (McLatchie et al., 2025).
To better understand the impact of this phenomenon on prediction, it is useful to distinguish between two regimes: when a statistical model is well-specified so that the data-generating process is recoverable by a singular parameter value, posterior concentration is the desired outcome: as more data accumulates, uncertainty contracts around the true parameter, and the posterior predictive distribution is well-calibrated. In practice however, statistical analysis is almost always subject to some degree of model misspecification. In this regime, no single parameter value is capable of describing the data generating process, so that the posterior predictives based on standard or generalised Bayesian posteriors become poorly calibrated (Syring and Martin, 2019). Importantly, this issue is not caused by obtaining predictive distributions through the marginalisation of a posterior, but by marginalising over a posterior which collapses to a point mass. This in turn provides strong motivation for methods whose parameter uncertainty can be guaranteed not to collapse to a point mass when this
@@
In this paper, we synthesise the above discussion into a new statistical principle for inference that results in a predictively oriented (PrO) posterior. The resulting approach is distinct from existing inference frameworks in two main ways: first, it treats prediction as its main goal without requiring a black box approach, while naturally inducing uncertainty quantification for a parameter of interest as a secondary inferential outcome. Second, it is adaptive: the PrO posterior collapses to a point mass only in the appropriate setting where the model is well-specified. When the model is misspecified however, PrO posteriors acknowledge that no single parameter is sufficient for capturing the true data-generating process, and yield an irreducible level of uncertainty, even with increasing sample size.
@@
Conceptually, PrO posteriors represent a useful intermediary between existing statistical paradigms and compromise between the two tenants of data modelling and algorithmic modelling sub-cultures in statistics that were famously identified by Breiman (2001). In practical terms, this means that PrO posteriors reconcile parameter inference and predictive inference in situations where the choice between either is difficult. Intriguingly, they do so through a posterior that expresses epistemological uncertainty in a way that is meaningfully distinct from standard and generalised Bayesian approaches.
@@
We strengthen this intuitive appeal by demonstrating that PrO posteriors possess strong theoretical guarantees and an algorithmically viable approach for computation. In particular, we formally establish that PrO posteriors predictively dominate classical and generalised Bayesian posteriors. Further, we characterise the type of parameter uncertainty they confer, and show that PrO posteriors are adaptive to model misspecification. Lastly, we introduce a sampling algorithm which can be used to sample from any such posterior, and illustrate its performance across a range of examples.
In the remainder, we motivate the need for PrO posteriors in more detail (Section 1), develop new mathematical tools for our theoretical analysis (Section 2), and then present our main findings (Section 3). A full discussion of all results and the assumptions we use is deferred to the end of the paper (Section 7). We then introduce computationally feasible algorithms for PrO posteriors (Section 4) and by illustrating their theoretical performance on a number of numerical examples (Section 5) before concluding with a detailed discussion of our findings (Section 6).
1 Motivation
@@
Throughout, we consider data ${x}_{1:n}=(x_{1},\dots,x_{n})$ for $x_{i}\in\mathcal{X}$ and $i=1,\dots,n$. We assume that each $x_{i}$ was generated independently from an unknown distribution $P_{0}$, and denote expectations of a function $f:\mathcal{X}\to\mathbb{R}$ with respect to $P_{0}$ by $\mathbb{E}[f(X)]$ or by $\mathbb{E}_{X\sim P_{0}}[f(X)]$ if we wish to emphasise the sampling distribution. To learn about $P_{0}$ from $x_{1:n}$, we follow the tenants of classical Bayesian analysis, and posit a model class $\mathcal{M}_{\Theta}:=\{P_{\theta}:\theta\in\Theta\}$ and a prior $\Pi\in\mathcal{P}(\Theta)$ over elements of $\mathcal{M}_{\Theta}$, where $\mathcal{P}(\Theta)$ denotes the space of probability measures on $\Theta$.
Instead of attempting to learn the best possible parameter, as in classical (or even generalised) Bayesian inference, we focus on learning the best possible predictive distribution. To rigorously formulate this optimality, we use scoring rules.
@@
Scoring rules are foundational statistical objects concerned with measuring accuracy of predictive distributions (see, e.g., Gneiting and Raftery, 2007). Given a target distribution $P_{0}$, a candidate predictive distribution $P$ is judged to be more accurate than an alternative candidate $P^{\prime}$ by an expected score $\mathcal{S}$ if and only if $\mathcal{S}(P,P_{0})\leq\mathcal{S}(P^{\prime},P_{0})$. The logarithmic scoring rule, given by $S(P,x)=-\log P(x)$, is the most common example and is related to the Kullback-Leibler divergence via $\mathrm{KL}(P^{\prime}\|P)=\mathcal{D}_{S}(P,P^{\prime})$. Other popular choices include the spherical score, cumulative ranked probability score (CRPS), the energy score, and squared maximum mean discrepancy (MMD), which for a kernel function $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{+}$ and the Dirac measure at $x$ denoted by $\updelta_{x}$ is given by
                                                                                                    $\mathrm{MMD}^{2}(P,\updelta_{x})$ ${}:={}$ $\mathbb{E}_{X\sim P,X^{\prime}\sim P}[k(X,X^{\prime})]-2\mathbb{E}_{X\sim P}[k(X,x)].$                                                                                                      (1)
Throughout the paper, we restrict our analysis to any score that satisfies the following minimal convexity assumption, which is satisfied for all scoring rules mentioned above.
Assumption 1.
@@ 1.1 Key Ideas
When $P_{\theta}$ has density $p_{\theta}$, the Bayes posterior $\Pi_{n}$ and its associated posterior predictive distribution are
                                                                                                    $\Pi_{n}(A):=\frac{\int_{A}p_{\theta}(x_{1:n})\mathsf{d}\Pi(\theta)}{\int_{\Theta}p_{\theta}(x_{1:n})\mathsf{d}\Pi(\theta)},\quad\quad P_{\Pi_{n}}:=\int_{\Theta}P_{\theta}\mathsf{d}\Pi_{n}(\theta).$                                                                                                      (2)
@@
The key motivation for this novel methodology is that while $P_{\Pi_{n}}$ is predictively optimal if $P_{0}\in\mathcal{M}_{\Theta}$ (see Aitchison, 1975), this optimality breaks down once $P_{0}\notin\mathcal{M}_{\Theta}$. Indeed, the Bayes posterior predictive can yield extremely poor predictive performance under even seemingly innocuous model misspecification (see, e.g., Grünwald and van Ommen, 2017). Though this is well-understood, robust Bayesian methods usually still treat prediction as a second-order concern, and are more concerned with parameter uncertainty. One family emblematic of this kind of research are Gibbs posteriors. (e.g. Jiang and Tanner, 2008; Bissiri et al., 2016; Zhang, 2006; Alquier, 2024; Syring and Martin, 2019; Martin and Syring, 2022; Knoblauch et al., 2018; Schmon et al., 2020; Chérief-Abdellatif and Alquier, 2020; Matsubara et al., 2022, 2023; Altamirano et al., 2023a, b; McLatchie et al., 2025; Duran-Martin et al., 2024; Laplante et al., 2025; Frazier et al., 2025). Many of these posteriors are based on an scoring rule evaluated on the sample, and given by, for $A\subseteq\Theta$,
@@
What then should we do if we wish to predict from a parametric model, but also suspect that $P_{0}\notin\mathcal{M}_{\Theta}$? Now, we argue that the quantity of interest should be the predictive $P_{Q}:=\int P_{\theta}\mathsf{d}Q(\theta)$, and should thus be scored directly. Doing so gives rise to the PrO posterior given by
@@
${Q}_{n}=\operatornamewithlimits{argmin}_{Q\in\mathcal{P}(\Theta)}\left\{\frac{\lambda_{n}}{n}\sum_{i=1}^{n}\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)+\mathrm{KL}(Q\|\Pi)\right\},$
where we write $\widehat{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)$ instead of ${S}\left(\int P_{\theta}\mathsf{d}Q(\theta),x_{i}\right)$ to highlight that throughout our theoretical and methodological development, PrO posteriors can be computed based on either exact or approximate predictive scoring rules.
1.2 Initial Comparison: Gibbs Posteriors vs PrO Posteriors
As we show later, PrO posteriors have several important properties, but the most fundamentally useful property is that they are adaptive: in contrast to Bayes and Gibbs posteriors, PrO posteriors adapt to model misspecification. We illustrate what this means by comparing the predictive distributions associated to $Q_{n}$ and $Q_{n}^{\dagger}$ under $\mathcal{M}_{\Theta}=\{P_{\theta}:\,\mathsf{N}(\theta,\sigma^{2}=1)\}$ for a range of data distributions. This model class is particularly educational here as the predictive is obtained from the posterior via smoothing with a Gaussian kernel with unit scaling. For predictive scores based on squared MMD, Figure 1 plots the results for the well-specified setting where $P_{0}\in\mathcal{M}_{\Theta}$ and three misspecified settings where $P_{0}\notin\mathcal{M}_{\Theta}$. While the PrO posterior behaves similarly to the MMD-based Gibbs posterior if the model is well-specified, it adapts to misspecification, and continues to produce reasonable predictive distributions across each setting. In contrast, Gibbs posteriors are blind to the mismatch between model and data.
As our example uses Gaussian likelihoods, one may misinterpret the phenomenon displayed as a cherry-picked pathology. In parametric models, however, Bayes posteriors and Gibbs posteriors are usually asymptotically Gaussian, and will concentrate around a single element of $\mathcal{M}_{\Theta}$. As we formally show later on, this property is not shared by PrO posteriors. On the contrary, depending on the severity of misspecification, they can be multi-modal and need not concentrate around a single element of $\mathcal{M}_{\Theta}$. Instead, they stabilise towards the predictively optimal average of elements in $\mathcal{M}_{\Theta}$ for $P_{0}$. As a result—and unlike Bayes posteriors and Gibbs posteriors—PrO posteriors produce predictively meaningful parameter uncertainty even under severe forms of misspecification.
Figure 1: Comparison of PrO posteriors and Gibbs posteriors in a well-specified and three misspecified regimes. Grey bars indicate a histogram of the observed data.
1.3 PrO Posteriors: A Third Modelling Culture?
In his prescient contribution, Breiman (2001) identified two divergent sub-cultures within statistical science. On the one hand, he spoke of a data modelling culture interested in hand-tailoring interpretable models for a given research problem and thus strongly associated to classical statistical concerns. On the other hand, he identified an increasing interest in machine learning and the associated algorithmic modelling culture focused on predictive accuracy and black box density estimation techniques. Does their reliance on a parametric base model $\mathcal{M}_{\Theta}$ make PrO posteriors part of the data modelling culture? Or does their focus on predictive performance make them into an offshoot of machine learning and the associated algorithmic modelling culture?
We believe that PrO posteriors constitute a meaningful and novel attempt to incorporate the most useful aspects of both modelling cultures: they rely on interpretable parameters born out of the data modelling culture, but gear inference towards predictive accuracy. When both interpretability and predictive performance are relevant, practitioners are often forced into a fraught choice: should they prefer an interpretable but predictively suboptimal model; or should they resort to a black box model that will improve predictive performance at the cost of interpretability and meaningful guarantees? PrO posteriors do not force us into this problematic binary choice: they rely on interpretable parameters born out of the data modelling culture, but gear inference towards predictive accuracy, and as a result perform well on both tasks. In Figure 2, we illustrate this by studying body measurements for different penguin species in the Palmer Archipelago (Horst et al., 2020). The PrO posterior not only identifies the data set’s fundamental structure by correctly discovering the three different species in it, but also produces a predictive distribution of excellent quality.
As we shall see in the remainder, this behaviour is not an accident: PrO posteriors provide an impactful third way of learning from data that combines predictive accuracy with interpretable uncertainty estimates. This behaviour is desirable for a wide array of scientific problems, including in digital twins, simulation models, causal modelling with observational data in public health and medicine, and forecasting applications in energy markets and finance. In operationalising a compromise between parameter inference and predictive accuracy, PrO posteriors also yield an altogether new kind of uncertainty quantification different from that of classical approaches. The next section elaborates on this, and explains why we can think of PrO posteriors as providing a novel type of epistemological uncertainty beyond what is currently attainable in classical Bayesian inference.
1.4 The Nature of PrO Posterior Uncertainty
To start, it is worth contrasting PrO posterior uncertainty expressed with a set of ideas and methods known under the umbrella of predictive Bayes (see e.g. Berti et al., 2021; Fortini and Petrone, 2020, 2023; Fong et al., 2023; Fortini and Petrone, 2025). Despite similar sounding names, the roles played by prediction in the construction of PrO posteriors and predictive Bayes approaches are fundamentally different: the PrO posterior is geared towards learning the predictive from seen data, predictive Bayes approaches instead posits a sequence of predictive distributions to model the unseen data. Put differently, while PrO posteriors make the predictive distribution the object of posterior inference, predictive Bayes methods specify a predictive distribution instead of a prior and likelihood.
In fact, specifying priors and likelihoods is just a convenient way of specifying a sequence of predictive distributions on unseen data: formalising this in a recent read paper at the Royal Statistical Society, Fong et al. (2023) used Doob’s Theorem to argue that the ultimate source of Bayesian uncertainty is unseen data. Taking this to its logical extreme, if we observe infinitely many data points, no uncertainty about the data remains and we can be perfectly confident in the conclusions we draw from data. This is in fact true more generally, as similar reasoning can be applied to various generalisations of Bayesian uncertainty, including the uncertainty expressed by Gibbs posteriors. One of the key implications is that both Bayesian and generalised Bayesian inference schemes result in a form of epistemological uncertainty that vanishes as the number of observations increase.
But should epistemological uncertainty indeed vanish as more data are observed? Though it may be tempting to answer this question in the affirmative, doing so would overlook an important practical limitation: for interpretability and computational reasons, statistical methods express epistemological uncertainty relative to a parameter $\theta\in\Theta$. Crucially, this implies that the only setting in which larger sample sizes should lead to a decrease in epistemological uncertainty is the rare situation where a singular value of $\theta\in\Theta$ can summarise all properties of $P_{0}$ that we are interested in. In other words, the kind of epistemological uncertainty conferred by (generalised) Bayesian methods must always be viewed relative only to the model class, not with respect to the choice of model class itself, rendering inference and particularly uncertainty quantification unreliable whenever $P_{0}\notin\mathcal{M}_{\Theta}$.
The epistemological uncertainty produced by PrO posteriors is more expansive: it delivers both parameter uncertainty and uncertainty about the inherent ability of $\mathcal{M}_{\Theta}$ to adequately explain the observed data. Indeed, when the model is misspecified, PrO posteriors generally do not collapse to a single element of $\mathcal{M}_{\Theta}$ as more data are observed. When $P_{0}\notin\mathcal{M}_{\Theta}$, they instead stabilise towards an irreducible form of uncertainty that originates in the fact that no single value of $\theta\in\Theta$ is sufficient to summarise all properties of $P_{0}$ we take an interest in. This is epistemological uncertainty in its truest form: because the model is misspecified, we should be uncertain about its ability to represent the observed data, even if we had observed the entire population $P_{0}$.
But what if the model is well-specified, so that $P_{0}\in\mathcal{M}_{\Theta}$? In this case, the epistemological uncertainty produced by Bayes and Gibbs posteriors is sufficient: there exists some $\theta\in\Theta$ that allows us to completely describe the observed data—and as more data accumulates, we should become increasingly certain about the ability of this value to correctly explain the data, and hence $P_{0}$. As it turns out, PrO posteriors square the circle: in the special case where the model is well-specified, they do behave like Bayes and Gibbs posteriors, and concentrate towards the parameter value that recovers $P_{0}$.
1.5 Contributions
We make three main contributions to demonstrate the advantages of PrO posteriors:
  1. C1.
    In Section 2, we establish a new mathematical representation that encompasses all PrO posteriors, as well as previous attempts at specifying similar objects.
  2. C2.
    In Sections 3 and 7, we summarise our substantive theoretical contributions. We formally establish that PrO posteriors predictively dominate Bayes posteriors and Gibbs posteriors alike, and prove that they are adaptive to model misspecification.
  3. C3.
    To make our theoretical contributions actionable, we also provide an asymptotically exact sampling strategy for PrO posteriors based on particle systems derived from Wasserstein gradient flows, and demonstrate its performance on several examples.
Our findings in C1 clarify and unify previous approaches, and craft a common lens through which to view a collection of somewhat disparate contributions over the years that used different justifications for objects that resemble PrO posteriors (e.g. Sheth and Khardon, 2020; Masegosa, 2020; Wei et al., 2021; Jankowiak et al., 2020b, a; Morningstar et al., 2022; Lai and Yao, 2024; Shen et al., 2025). Beyond that, C1 powers much of the theory we develop in C2, which in turn provides a rigorous justification for heuristic strategies in earlier literature which targeted variational approximations of $Q_{n}$ for the special case of the log score (see Sheth and Khardon, 2020; Wei et al., 2021; Jankowiak et al., 2020b, a). One important part of this new theory is a significantly strengthened and generalised version of the PAC-Bayes type bounds in related approaches, such as in Masegosa (2020) and Morningstar et al. (2022) and Morningstar et al. (2022) in Lemma 1. Unlike these previous attempts, our result is not restricted to any one particular score or approximation, but instead applies universally under mild regularity conditions. Despite the generality of our approach, we manage to dramatically improve previous bounds for the special case of the logarithmic scoring rule. In fact, our bounds are sufficiently sharp to allow for a characterisation of the qualitative behaviours of PrO posteriors (see Theorems 12). Lastly, C3 provides a generic approach for drawing samples from any PrO posterior—despite the fact that such objects generally have no known analytic form. This is very significant: aside from the special case of MMD-based PrO posteriors in Shen et al. (2025), all previous work had been focused on variational approximations (see e.g. Sheth and Khardon, 2020; Jankowiak et al., 2020b, a; Lai and Yao, 2024), and it was not clear whether a PrO posterior constructed from any other score could even be sampled from at all. As we show, this can indeed be achieved by suitably building on the Wasserstein gradient flow and the ideas of Wild et al. (2023), Wild et al. (2025), and Shen et al. (2025).
2 Background and Theoretical Groundwork
To set the stage for our theoretical developments, we introduce three key ideas that guide our methodological and theoretical innovations. In Section 2.1, we study representations and approximations of scoring rules whose structure leads to a tractable objective, from which we derive a novel PAC-Bayes type bound in Section 2.2, which serves as the foundation of our investigation into the predictive performance of PrO posteriors.
2.1 Exact and Approximate Representations of Predictive Scores
To simplify language and to highlight that PrO posteriors apply scoring rules to the predictive distribution $P_{Q}=\int P_{\theta}\mathsf{d}Q(\theta)$, the remainder of this paper will use the shorthand predictive scoring rule to refer to a scoring rule which is applied to the predictive distribution $P_{Q}$. One of the challenges of working with predictive scoring rules is that they do not admit simple representations. However, Assumption 1 and Jensen’s inequality imply that
                                                                                                    $S\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)$ ${}={}$ $\int S\left(P_{\theta},x\right)\mathsf{d}Q(\theta)-\Delta(Q,x),$                                                                                                      (5)
for all $x\in\mathcal{X}$ and some non-negative remainder term $\Delta:\mathcal{P}(\Theta)\times\mathcal{X}\to\mathbb{R}_{+}$. This representation has important practical ramifications: depending on the form of $\Delta(Q,x)$, a predictive score is either intractable, or can be tractably approximated. In the remainder, we call a predictive scoring rule tractable when it can be written as the expectation over a product measure for any $x\in\mathcal{X}$; so that for some $k\in\mathbb{N}_{+}$ and a loss $L:\Theta^{k}\times\mathcal{X}\rightarrow\mathbb{R}$, one obtains
$S\left(\int P_{\theta}\mathsf{d}Q(\theta),x\right)=\int L(\theta_{1:k},x)\mathsf{d}Q^{k}(\theta_{1:k}),\quad Q^{k}(\theta_{1:k})=\bigotimes_{i=1}^{k}Q(\theta_{i}).$ (6)
If one can sample from $Q$, this allows approximation to arbitrary precision via Monte Carlo methods. For example, kernel scores, such as MMD and CRPS, are tractable with $k=2$ (see Lemma 3 in Appendix B).
While tractable predictive scoring rules facilitate computation, several predictive scores of interest are intractable—most notably the logarithmic predictive scoring rule (see Lemma 4 in Appendix B). From the decomposition in (2.1), its associated Jensen’s gap depends on an intermediate value between $p_{\theta}(x)$ and $\int p_{\theta}(x)\mathsf{d}Q(\theta)$ denoted $g_{x}(p_{\theta},Q)$, and is

### v2

@@ Abstract
We advocate for a new statistical principle that combines the most desirable aspects of both parameter inference and density estimation. This leads us to the predictively oriented (PrO) posterior, which expresses uncertainty as a consequence of predictive ability. Doing so leads to inferences which predictively dominate both classical and generalised Bayes posterior predictive distributions: up to logarithmic factors, PrO posteriors converge to the predictively optimal model average. Whereas classical and generalised Bayes posteriors only achieve this rate if the model can recover the data-generating process, PrO posteriors adapt to the level of model misspecification. This means that they concentrate around the true model in the same way as Bayes and Gibbs posteriors if the model can recover the data-generating distribution, but do not concentrate in the presence of non-trivial forms of model misspecification. Instead, they stabilise towards a predictively optimal posterior whose degree of irreducible uncertainty admits an interpretation as the degree of model misspecification—a sharp contrast to how Bayesian uncertainty and its existing extensions behave. Lastly, we show that PrO posteriors can be sampled from by evolving particles based on mean field Langevin dynamics, and verify the practical significance of our theoretical developments on a number of numerical examples.
@@
In the context of prediction however, posterior collapse is generally not desirable: it implies that as the sample size increases, marginalising over the Bayes posterior ceases to function as a genuine combination of individual predictive distributions. Instead of averaging over models, such predictive distributions essentially predict from a single model. Crucially, this behaviour persists even for generalised Bayesian methods that replace a likelihood-based approach with a loss-based belief update (Bissiri et al., 2016; Jiang and Tanner, 2008; Zhang, 2006; Knoblauch et al., 2022; Alquier, 2024; Syring and Martin, 2019; Martin and Syring, 2022). As a result, the posterior predictives obtained from both classical and generalised Bayes approaches quickly resemble those obtained from a point estimator as sample size grows (McLatchie et al., 2025).
To better understand the impact of this phenomenon on prediction, it is useful to distinguish between two regimes. When a statistical model is well-specified so that the data-generating process is recoverable by a singular parameter value, posterior concentration is the desired outcome: as more data accumulates, uncertainty contracts around the true parameter, and the posterior predictive distribution is well-calibrated. In practice however, statistical analysis is almost always subject to some degree of model misspecification. In this regime, no single parameter value is capable of describing the data generating process, so that the posterior predictives based on standard or generalised Bayesian posteriors become poorly calibrated (Syring and Martin, 2019). Importantly, this issue is not caused by obtaining predictive distributions through the marginalisation of a posterior, but by marginalising over a posterior which collapses to a point mass. This in turn provides strong motivation for methods whose parameter uncertainty can be guaranteed not to collapse to a point mass when this
@@
In this paper, we synthesise the above discussion into a new statistical principle for inference that results in a predictively oriented (PrO) posterior. The resulting approach is distinct from existing inference frameworks in two main ways: first, it treats prediction as its main goal without requiring a black box approach, while naturally inducing uncertainty quantification for a parameter of interest as a secondary inferential outcome; second, it is adaptive: the PrO posterior collapses to a point mass only in the appropriate setting where the model is well-specified. When the model is misspecified however, PrO posteriors acknowledge that no single parameter is sufficient for capturing the true data-generating process, and yield an irreducible level of uncertainty, even with increasing sample size.
@@
Conceptually, PrO posteriors represent a useful intermediary between existing statistical paradigms and compromise between the two tenants of data modelling and algorithmic modelling sub-cultures in statistics that were famously identified by Breiman (2001). In practical terms, this means that PrO posteriors reconcile parameter inference and predictive inference in situations where the choice between either is difficult. Intriguingly, they do so through a posterior that expresses epistemic uncertainty in a way that is meaningfully distinct from standard and generalised Bayesian approaches.
@@
We strengthen this intuitive appeal by demonstrating that PrO posteriors possess strong theoretical guarantees and an algorithmically viable approach for computation. To do so, we first set the scene and formally introduce PrO posteriors (Section 1). Next, we investigate the nature of the uncertainty which they confer, show that they are adaptive to model misspecification, and explain their broader place in the statistical sciences (Section 2). From there, we go on to formally establish that PrO posteriors predictively dominate classical and generalised Bayesian posteriors, and characterise their asymptotic behaviour (Sections 3 and 8). We end by introducing a sampling algorithm which can be used to sample from PrO posteriors (Section 6), exploring their empirical performance on a range of examples (Section 5), and providing a discussion of our findings (Section 7).
1 Background
@@
Throughout, we consider data ${x}_{1:n}=(x_{1},\dots,x_{n})$ for $x_{i}\in\mathcal{X}$ and $i=1,\dots,n$. We assume that ${x}_{1:n}$ was generated from an unknown distribution $P_{0}$. With this, we denote expectations of a function $f:\mathcal{X}\to\mathbb{R}$ with respect to $P_{0}$ by $\mathbb{E}[f(X)]$ or by $\mathbb{E}_{X\sim P_{0}}[f(X)]$ if we wish to emphasise the sampling distribution. To learn about $P_{0}$ from $x_{1:n}$, we follow the tenants of classical Bayesian analysis, and posit a model class $\mathcal{M}_{\Theta}:=\{P_{\theta}:\theta\in\Theta\}$ and a prior $\Pi\in\mathcal{P}(\Theta)$ over elements of $\mathcal{M}_{\Theta}$, where $\mathcal{P}(\Theta)$ denotes the space of probability measures on $\Theta$. Instead of attempting to learn the best possible parameter, as in classical (or even generalised) Bayesian inference, we focus on learning the best possible predictive distribution. To rigorously formulate this optimality, we use scoring rules.
@@
Scoring rules are foundational statistical objects concerned with measuring accuracy of predictive distributions (see, e.g., Gneiting and Raftery, 2007). Given a target distribution $P_{0}$, a candidate predictive distribution $P$ is judged to be more accurate than an alternative candidate $P^{\prime}$ by an expected score $\mathcal{S}$ if and only if $\mathcal{S}(P,P_{0})\leq\mathcal{S}(P^{\prime},P_{0})$. The logarithmic scoring rule, given by $S(P,x)=-\log\mathsf{d}P(x)$, is the most common example and is related to the Kullback-Leibler divergence via $\mathrm{KL}(P^{\prime}\|P)=\mathcal{D}_{S}(P,P^{\prime})$. Other popular choices include the spherical score, cumulative ranked probability score (CRPS), the energy score, and squared maximum mean discrepancy (MMD), which for a kernel function $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{+}$ and the Dirac measure at $x$ denoted by $\updelta_{x}$ is given by $\mathrm{MMD}^{2}(P,\updelta_{x}):=\mathbb{E}_{X\sim P,X^{\prime}\sim P}[k(X,X^{\prime})]+k(x,x)-2\mathbb{E}_{X\sim P}[k(X,x)].$
@@ 1.1 Key Ideas
The Bayes posterior $\Pi_{n}$ and its associated posterior predictive distribution are
                                                                                                    $\Pi_{n}(A):=\frac{\int_{A}\mathsf{d}P_{\theta}(x_{1:n})\mathsf{d}\Pi(\theta)}{\int_{\Theta}\mathsf{d}P_{\theta}(x_{1:n})\mathsf{d}\Pi(\theta)},\quad\quad P_{\Pi_{n}}:=\int_{\Theta}P_{\theta}\mathsf{d}\Pi_{n}(\theta).$                                                                                                      (1)
@@
The key motivation for this novel methodology is that while $P_{\Pi_{n}}$ is predictively optimal if $P_{0}\in\mathcal{M}_{\Theta}$ (see Aitchison, 1975), this optimality breaks down once $P_{0}\notin\mathcal{M}_{\Theta}$. Indeed, the Bayes posterior predictive can yield extremely poor predictive performance under even seemingly innocuous model misspecification (see Grünwald and van Ommen, 2017). Though this is well-understood, robust Bayesian methods usually still treat prediction as a second-order concern, and are more concerned with parameter uncertainty. One family emblematic of this kind of research are Gibbs posteriors (e.g. Zhang, 2006; Jiang and Tanner, 2008; Jewson et al., 2018; Knoblauch et al., 2018; Giummolè et al., 2019; Syring and Martin, 2019; Schmon et al., 2020; Chérief-Abdellatif and Alquier, 2020; Matsubara et al., 2022; Martin and Syring, 2022; Matsubara et al., 2023; Altamirano et al., 2023a, b; Duran-Martin et al., 2024; McLatchie et al., 2025; Laplante et al., 2025; Frazier et al., 2025). Many of these posteriors are based on an scoring rule evaluated on the sample, and for $A\subseteq\Theta$ are given by
@@
What then should we do if we wish to conduct inference on model parameters, but wish to ensure that the prediction made under the model are accurate, while also suspecting that $P_{0}\notin\mathcal{M}_{\Theta}$? We argue that, while the parameters of interest remain $\theta\in\Theta$, we should ensure that the beliefs are generated in such a way that the resulting posterior predictive $P_{Q}:=\int P_{\theta}\mathsf{d}Q(\theta)$ is accurate. To do so, we must score the predictive $P_{Q}$ directly, which gives rise to what we shall call the predictively oriented (PrO) posterior given by
@@
${Q}_{n}=\operatornamewithlimits{argmin}_{Q\in\mathcal{P}(\Theta)}\left\{\frac{\lambda_{n}}{n}\sum_{i=1}^{n}S\left(P_{Q},x_{i}\right)+\mathrm{KL}(Q\|\Pi)\right\}.$ (4)
The PrO posterior can also be represented explicitly through a fixed point equation as